OpenStudy (shamil98):

A particle starts at x = 0 and moves along the x-axis at velocity v(t) = 5 for time t ≥ 0. Where is the particle at t = 4?

OpenStudy (agent47):

draw isn't working....

OpenStudy (agent47):

v(t) = 5 Distance is the integral of velocity from 0 to 5.

OpenStudy (perl):

position is the integral of velocity, so x(t) = 5*t + c , and since x(0) = 0, we have x(0) = 5*0 + c , c = 0. so x(t) = 5t , and at t= 4 x(4) = 5*4 = 20 units

OpenStudy (agent47):

And you don't have to worry about splitting it up as it's in one direction only.

OpenStudy (perl):

20 units in the positive direction (where you defined positive as right or left)

OpenStudy (shamil98):

so basically you integrate the equation v(t) = 5 and do what you did for questions like these?

OpenStudy (perl):

yes